OpenStudy (anonymous):

Carrots sell for $2.10 per pound, and crackers sell for $2.90 per pound. Glen bought some carrots and some crackers. The total weight was 2.3 pounds and cost $6.03. Let p represent the number of pounds of carrots. Which equation represents the situation described?

OpenStudy (anonymous):

well hello there death the kid....

OpenStudy (anonymous):

welcome to openstudy

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

mhmm. so do you have answer choices? the question seems like there are some...

OpenStudy (anonymous):

A. –0.8p + 6.03 = 6.67 B. 2.1p + 2.9 = 6.03 C. –0.8p + 6.67 = 6.03 D. 2.1p + 6.03 = 2.9

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

no, its okay, so p is carrots.... and the cost of carrots is $2.1 crackers is just a normal number 2.9 and it all totals up to 6.03 so you would have... 2.1p + 2.9 = 6.03

OpenStudy (anonymous):

i believe that is b

OpenStudy (anonymous):

i hope this helps @Death_The_Kid

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thank you SO much!!! :D

OpenStudy (anonymous):

anytime :D feel free to ask me if you need help :D

OpenStudy (anonymous):

Believe me I will

OpenStudy (anonymous):

:D